# Gravothermal Catastrophe in Resonant Self-interacting Dark Matter Models

### v1

@@ I Introduction
We consider a class of SIDM models in which a resonance in the cross section causes an order-of-magnitude enhancement to the scattering cross section near a particle relative velocity. Resonances refer to an enhancement or suppression of the cross section that results from the formation of bound states between DM particles and attractive potential. Resonances can manifest as single pronounced peaks, multiple peaks, or as suppression or enhancement of the cross section strength across a range of speeds, depending on the details of the particle physics model [9, 47, 39]. For this class of model, the cross section amplitude can signifcantly change with radius inside a halo, complicating the relatively simple physical picture in which a halo evolves under the influence of a single effective cross section. This class model of model has received relatively limited study. Ref. [48] examined the evolution of a halo analytically using the gravothermal fluid formalism, while Ref. [28] studied a resonant model in a numerical simulation, but did not distinguish its structural evolution from models with simpler velocity-dependence.
@@ II Particle physics model
We consider an SIDM cross section that results from interactions between DM particles of mass $m_{\chi}$ through a force carrier of mass $m_{\phi}$. We consider the case of an attractive Yukawa potential $\left(c=\hbar=1\right)$
@@
where $\alpha$ determines the strength of the potential. While no exact closed-form solutions exist for the differential scattering cross section corresponding to $V\left(r\right)$, in certain regimes of parameter space various approximations give analytic and closed-form solutions that match the exact cross section to high precision. For example, for weak potentials $\left(\alpha m_{\chi}/m_{\phi}<<1\right)$, one obtains an approximate analytic expression for the differential scattering cross section using the Born approximation
@@
where $v$ represents relative velocity, $\sigma_{0}\equiv\alpha^{2}m_{\chi}^{2}/m_{\phi}^{4}$ and $w\equiv m_{\phi}/m_{\chi}$.
@@
It has become common practice to average over the angular dependence of the differential scattering cross section when predicting the internal structure of DM halos in SIDM. This simplification can yield accurate predictions for the evolution of halo profiles provided one chooses a suitable proxy for ${\rm d}\sigma/{\rm d}\Omega$. As shown by Yang and Yu [20], the viscosity transfer cross section $\sigma_{V}=\int\frac{{\rm d}\sigma}{{\rm d}\Omega}\sin^{2}\theta{\rm d}\theta$, for which $\sigma_{V}$ is given by [49]
@@
For the analysis presented in this work, we consider the resonant cross section represented by the shaded region in the left-hand panel Figure 1. This cross section was also considered by Ref. [39] in a strong-lensing analysis of multiply-imaged quasars, and has $m_{\rm{\chi}}=38.1\ \rm{GeV}$, $m_{\rm{\phi}}=6.8\ \rm{MeV}$, and $\alpha=1.6\times 10^{-3}$. The amplitude of the cross section, as indicated by the axis labels on the right-hand side of the figure, reaches $\sim 100\ \rm{cm^{2}}\ \rm{g^{-1}}$ at $v\sim 16\ \rm{km}\ \rm{s^{-1}}$, a feature that arises primarily from constructive interference between the $\ell=3$ and $\ell=1$ partial waves. To explore the consequences of this model for structure formation, we run a N-body simulation of a DM halo evolving under the influence of this resonant cross section, and track its evolution from core formation until core collapse.
@@ VI Discussion and conclusions
A modified calculation of the core collapse timescale based on a time-dependent definition of the thermal conduction cross section $\sigma_{\kappa}$ provides an improved description of the collapse phase for the halo evolving with a resonant cross section. This suggests that a modified calculation of the thermal conduction cross section could restore universality for resonant cross sections, provided this modified treatment reverts to the definition of $\sigma_{\kappa}$ for velocity-independent cross sections. Exploring this topic further, and simulating the evolution of halos with different resonant cross sections, will be the topic of future study.
@@ Acknowledgements.
We thank Annika Peter, Antti Asikainen, and Yi-Ming Zhong for useful discussions. DG acknowledges support for this work provided by the Brinson Foundation through a Brinson Prize Fellowship grant. MV acknowledges support through NASA ATP 19-ATP19-0019, 19-ATP19-0020, 19-ATP19-0167, and NSF grants AST-1814053, AST-1814259, AST-1909831, AST-2007355 and AST-2107724.

### v2

@@ I Introduction
We consider a class of SIDM models in which a resonance in the cross section causes an order-of-magnitude enhancement to the scattering cross section near a particle relative velocity. Resonances refer to an enhancement or suppression of the cross section that results from the formation of bound states between DM particles and attractive potential. Resonances can manifest as single pronounced peaks, multiple peaks, or as suppression or enhancement of the cross section strength across a range of speeds, depending on the details of the particle physics model [9, 47, 39]. For this class of model, the cross section amplitude can significantly change with radius inside a halo, complicating the relatively simple physical picture in which a halo evolves under the influence of a single effective cross section. This class model of model has received relatively limited study. Ref. [48] examined the evolution of a halo analytically using the gravothermal fluid formalism, while Ref. [28] studied a resonant model in a numerical simulation, but did not distinguish its structural evolution from models with simpler velocity-dependence.
@@ II Particle physics model
We consider a SIDM cross section that results from interactions between DM particles of mass $m_{\chi}$ through a force carrier of mass $m_{\phi}$. We consider the case of an attractive Yukawa potential $\left(c=\hbar=1\right)$
@@
where $r$ is the distance between DM particles, and $\alpha$ determines the strength of the potential. While no exact closed-form solutions exist for the differential scattering cross section corresponding to $V\left(r\right)$, in certain regimes of parameter space various approximations give analytic and closed-form solutions that match the exact cross section to high precision. For example, for weak potentials $\left(\alpha m_{\chi}/m_{\phi}<<1\right)$, one obtains an approximate analytic expression for the differential scattering cross section using the Born approximation
@@
where $v$, $\theta$, and $\Omega$ represent relative velocity, scattering angle, and the solid angle, respectively. The parameters $\sigma_{0}$ and $w$ are related to the particle physics model by $\sigma_{0}\equiv\alpha^{2}m_{\chi}^{2}/m_{\phi}^{4}$ and $w\equiv m_{\phi}/m_{\chi}$.
@@
It has become common practice to average over the angular dependence of the differential scattering cross section when predicting the internal structure of DM halos in SIDM. This simplification can yield accurate predictions for the evolution of halo profiles provided one chooses a suitable proxy for ${\rm d}\sigma/{\rm d}\Omega$. As shown by Yang and Yu [20], the viscosity transfer cross section $\sigma_{V}=\int\frac{{\rm d}\sigma}{{\rm d}\Omega}\sin^{2}\theta{\rm d}\Omega$, for which $\sigma_{V}$ is given by [49]
@@
For the analysis presented in this work, we consider the resonant cross section represented by the shaded region in the left-hand panel Figure 1. This cross section was also considered by Ref. [39] in a strong-lensing analysis of multiply-imaged quasars, and has $m_{\rm{\chi}}=31.8\ \rm{GeV}$, $m_{\rm{\phi}}=5.7\ \rm{MeV}$, and $\alpha=1.6\times 10^{-3}$. The amplitude of the cross section, as indicated by the axis labels on the right-hand side of the figure, reaches $\sim 100\ \rm{cm^{2}}\ \rm{g^{-1}}$ at $v\sim 16\ \rm{km}\ \rm{s^{-1}}$, a feature that arises primarily from constructive interference between the $\ell=3$ and $\ell=1$ partial waves. To explore the consequences of this model for structure formation, we run a N-body simulation of a DM halo evolving under the influence of this resonant cross section, and track its evolution from core formation until core collapse.
@@ VI Discussion and conclusions
A modified calculation of the core collapse timescale based on a time-dependent definition of the thermal conduction cross section $\sigma_{\kappa}$ provides an improved description of the collapse phase for the halo evolving with a resonant cross section. This suggests that a modified calculation of the thermal conduction cross section could restore universality for resonant cross sections, provided this modified treatment reverts to the definition of $\sigma_{\kappa}$ for velocity-independent cross sections. Additionally, provided analyses are robust to systematic uncertainties at the level of $\sim 20\%$ in the collapse timescale, studies involving dwarf galaxy observations and gravitational lensing could distinguish these classes of models from collisionless dark matter, improving on the existing constraints from lensing on RCS models presented by [39]. Further exploration of halo evolution with resonant SIDM cross sections with different resonant structures across a wider range of halo masses will be presented in a forthcoming publication.
@@ Acknowledgements.
We thank Annika Peter, Antti Asikainen, and Yi-Ming Zhong for useful discussions. We thank the anonymous referee for a constructive report. DG acknowledges support for this work provided by the Brinson Foundation through a Brinson Prize Fellowship grant. MV acknowledges support through NASA ATP 19-ATP19-0019, 19-ATP19-0020, 19-ATP19-0167, and NSF grants AST-1814053, AST-1814259, AST-1909831, AST-2007355 and AST-2107724.